# How Transformers Implement Induction Heads: Approximation and Optimization Analysis

### v2

@@ 5.1.1 Mixed Target Function
$f_{\textnormal{{G}}_{4}}^{\star}(X):=x_{T-2},\quad f_{\textnormal{{IH}}_{2}}^{\star}(X):=\sum_{s=2}^{L-1}{\rm softmax}\left(x_{L}{w^{\star}}^{2}x_{s-1}\right)x_{s}.$
@@
Since the transformer studied in this section does not have FFNs, its expressive power is limited. Consequently, we only consider the simple but representative mixed target (8). However, (8) can be generalized to $f^{\star}(X)=F(f_{\textnormal{{G}}_{4}}^{\star}(X);f_{\textnormal{{IH}}_{2}}^{\star}(X))$, where $F$ is general nonlinear function. Such a form can be efficiently approximated by transformers with FFNs. We leave the optimization analysis under this general setting for future work.
@@ 5.1.2 Two-layer Multi-head Transformer with Reparameterization
    $y_{s}=\sum_{\tau=1}^{s-1}{\rm softmax}\left(-p^{(1,1)}(s-1-\tau)\right)x_{\tau},\ s\in[L],$ (9)
    where $p^{(1,1)}$, used in RPE (3), is the unique trainable parameter in this layer.
@@
    $W_{Q}^{(2,1)}=W_{K}^{(2,1)}=0,W_{V}^{(2,1)}=\begin{pmatrix}0&w_{V}^{(2,1)}\\ 0&0\end{pmatrix},p^{(2,2)}=0,W_{V}^{(2,2)}=\begin{pmatrix}w_{V}^{(2,2)}&0\\ 0&0\end{pmatrix}.$
@@
    ${\textnormal{{TF}}}_{-1}(X;\theta)=\begin{pmatrix}w_{V}^{(2,1)}\sum_{s=2}^{L-2}{\rm softmax}(-p^{(2,1)}(L-1-s))y_{s}\\ w_{V}^{(2,2)}\sum_{s=2}^{L-2}{\rm softmax}\left(x_{L}w_{Q}^{(2,2)}w_{K}^{(2,2)}x_{s-1}\right)x_{s}\end{pmatrix},$ (10)
@@ 5.1.3 Gradient Flow on Square Loss
$\mathcal{L}_{\textnormal{{G}}_{4}}(\theta)=\frac{1}{2}{\mathbb{E}}_{X}\left({\textnormal{{TF}}}_{-1,1}({X};\theta)-f_{1}^{\star}(X)\right)^{2},\ \mathcal{L}_{\textnormal{{IH}}_{2}}(\theta)=\frac{1}{2}{\mathbb{E}}_{X}\left({\textnormal{{TF}}}_{-1,2}({X};\theta)-f_{2}^{\star}(X)\right)^{2},$
@@
$\frac{{\textnormal{d}}\theta(t)}{{\textnormal{d}}t}=-\nabla\mathcal{L}(\theta(t)),\text{\ starting with\ }\theta(0)=(\sigma_{\rm init},\cdots,\sigma_{\rm init})^{\top},$ (12)
@@
According to (9), this lemma implies that, at the end of Training Stage I, the first layer captures the preceding token $x_{s-1}$ for each token $x_{s}$, i.e., $y_{s}=x_{s-1}$. This property is crucial for transformers to implement induction heads and aligns with our approximation result in Theorem 4.1. The proof of Lemma 5.3 is deferred to Appendix B.
@@ 5.2 Training Stage II: Transition from $4$-gram to Induction Head
Lemma 5.4 is similar to the balance result for homogeneous networks (Du et al., 2018), and its proof can be found at the start of Appendix C. By this lemma, we can define $w_{KQ}^{(2,2)}:=w_{Q}\equiv w_{K}$. Additionally, Lemma 5.3 ensures that $p^{(1,1)}=+\infty$ holds during Stage II. For simplicity, we denote $w_{V_{1}}:=w_{V}^{(2,1)},w_{V_{2}}:=w_{V}^{(2,2)},p:=p^{(2,1)},w_{KQ}:=w_{KQ}^{(2,2)}$. Consequently, the training dynamics are reduced for four parameters
@@
where we still denote the set of parameters as $\theta$ without loss of ambiguity. Note that the problem remains highly non-convex due to the joint optimization of both inner parameters ($p,w_{KQ}$) and outer parameters ($w_{V_{1}},w_{V_{2}}$) in the two heads. In this stage, GF has a unique fixed point: $w_{V_{1}}=\frac{\alpha^{\star}}{1+\alpha^{\star}},w_{V_{2}}=\frac{1}{1+\alpha^{\star}},{p}=+\infty,w_{KQ}=w^{\star}$, which corresponds to a global minimizer of the objective (11).
Figure 2: Visualization of the total loss, partial loss, and the parameter dynamics in our setting with $\alpha^{\star}=1,w^{\star}=0.49,\sigma_{\rm init}=0.01,L=1000$. The figure clearly shows that transformer learns the $4$-gram component first and then, starts to learn the induction head mechanism. Notably, the entire dynamics exhibit four phases.
As shown in Figure 2, a learning transition from the 4-gram mechanism to the induction head mechanism does occur in our setting. Moreover, the learning process exhibits a four-phase dynamics. The next theorem provides a precise characterization of the four phases, whose proof can be found in Appendix C.
@@
    $\mathcal{L}_{\textnormal{{G}}_{4}}(\theta(T_{\rm I}))\leq 0.01\cdot\mathcal{L}_{\textnormal{{G}}_{4}}(\theta(0)),\quad\mathcal{L}_{\textnormal{{IH}}_{2}}(\theta(T_{\rm I}))\geq 0.99\cdot\mathcal{L}_{\textnormal{{IH}}_{2}}(\theta(0)).$
@@
    $T_{\rm II}$ $:=\inf\Big{\{}t>T_{o}:\mathcal{L}_{\textnormal{{IH}}_{2}}(\theta(t))\leq 0.99\cdot\mathcal{L}_{\textnormal{{IH}}_{2}}(\theta(T_{o}))\Big{\}}=\Theta\left((\alpha^{\star}+1)^{2}L\log(1/\sigma_{\rm init})/{w^{\star}}^{2}\right);$
    $T_{\rm III}$ $:=\inf\Big{\{}t>T_{o}:\mathcal{L}_{\textnormal{{IH}}_{2}}(\theta(t))\leq 0.01\cdot\mathcal{L}_{\textnormal{{IH}}_{2}}(\theta(T_{o}))\Big{\}}=\Theta\left((\alpha^{\star}+1)^{2}L\log(1/\sigma_{\rm init})/{w^{\star}}^{2}\right).$
@@
    $w_{KQ}(t)=\sigma_{{\rm init}}\cdot\exp\left(\Theta\left(\frac{{w^{\star}}^{2}t}{(1+\alpha^{\star})^{2}L}\right)\right),\ t<T_{\rm III}.$
  • Phase IV (convergence). In this phase, the loss convergences. Specifically, the following convergence rates hold for all $t>T_{\rm III}:$
    ${\mathcal{L}_{\textnormal{{G}}_{4}}(\theta(t))}=\mathcal{O}\left(\frac{1}{t}\right),\quad\mathcal{L}_{\textnormal{{IH}}_{2}}(\theta(t))=\mathcal{O}\left(\exp\left(-\Omega\left(\frac{{w^{\star}}^{2}t}{(1+\alpha^{\star})^{2}L}\right)\right)\right),$
@@
Proof idea. Our analysis relies on a precise description of the parameter trajectories, driven by two key observations: 1) the dynamics of two heads can be decoupled; 2) there exist a distinct transition point in the dynamics of each head, as shown in Figure 2 (right). These insights lead us to divide the analysis of each head into two phases: a monotonic phase and a convergence phase. Particularly, for the convergence phase, we introduce a novel Lyapunov function that leverages the unique dynamical structure of self-attention. This Lyapunov function may be of independent interest and offers potential for studying broader issues in self-attention dynamics.
@@ 6 Conclusion
Appendix A Proofs in Section 4
A.1 Proof of Theorem 4.1
$\textnormal{{IH}}_{2}(X_{L})=\sum_{s=2}^{L-1}{\rm softmax}\big{(}x_{L}^{\top}W^{\star}x_{s-1}\big{)}x_{s},$ (13)
Theorem A.1 (Restatement of Theorem 4.1).

### v3

@@ 4.1.1 Mixed Target Function
$f_{\textnormal{{G}}_{4}}^{\star}(X)$ $:=x_{L-2},$
$f_{\textnormal{{IH}}_{2}}^{\star}(X)$ $:=\sum_{s=2}^{L-1}x_{s}\ {\rm sm}\Big{(}\big{(}x_{L}{w^{\star}}^{2}x_{\nu-1}\big{)}_{\nu=2}^{L-1}\Big{)}_{\nu=s}.$
@@
Since the transformer studied in this section does not have FFNs, its expressive power is limited. Consequently, we only consider the simple but representative mixed target (8). However, (8) can be generalized to $f^{\star}(X)=F(f_{\textnormal{{G}}_{4}}^{\star}(X);f_{\textnormal{{IH}}_{2}}^{\star}(X))$, where $F$ is general nonlinear function. Such a form can be efficiently approximated by transformers with FFNs. We leave the dynamics analysis under this general setting for future work.
@@ 4.1.2 Two-layer Multi-head Transformer with Reparameterization
    $y_{s}\!=\!\sum_{\tau=1}^{s-1}x_{\tau}\ {\rm sm}\left(\!\big{(}-p^{(1,1)}(s-1-\nu)\big{)}_{\nu=1}^{s-1}\right)_{\nu=\tau}$ (9)
    for $s\in[L]$, where $p^{(1,1)}$, used in RPE (3), is the unique trainable parameter in this layer.
@@
    $W_{Q}^{(2,1)}=W_{K}^{(2,1)}=0,W_{V}^{(2,1)}=\begin{pmatrix}0&w_{V}^{(2,1)}\\ 0&0\end{pmatrix},$
    $p^{(2,2)}=0,W_{V}^{(2,2)}=\begin{pmatrix}w_{V}^{(2,2)}&0\\ 0&0\end{pmatrix}.$
@@
    $\!\!{\textnormal{{TF}}}_{-1}(X;\theta)\!=\!\Big{(}\sum_{s=2}^{L-2}w_{V}^{(2,1)}y_{s}\pi_{s},\sum_{s=2}^{L-2}w_{V}^{(2,2)}x_{s}\rho_{s}\!\Big{)}^{\top}\!\!,$
    $\begin{gathered}\pi_{s}={\rm sm}\left(\big{(}-p^{(2,1)}(L-1-\nu)\big{)}_{\nu=2}^{L-2}\right)_{\nu=s},\\ \rho_{s}={\rm sm}\left(\big{(}x_{L}w_{Q}^{(2,2)}w_{K}^{(2,2)}x_{\nu-1}\big{)}_{\nu=2}^{L-2}\right)_{\nu=s},\end{gathered}$ (10)
@@ 4.1.3 Gradient Flow on Square Loss
$\mathcal{L}_{\textnormal{{G}}_{4}}(\theta)$ $=\frac{1}{2}{\mathbb{E}}_{X}\left({\textnormal{{TF}}}_{-1,1}({X};\theta)-f_{1}^{\star}(X)\right)^{2},$
$\mathcal{L}_{\textnormal{{IH}}_{2}}(\theta)$ $=\frac{1}{2}{\mathbb{E}}_{X}\left({\textnormal{{TF}}}_{-1,2}({X};\theta)-f_{2}^{\star}(X)\right)^{2},$
@@
$\begin{gathered}\frac{{\textnormal{d}}\theta(t)}{{\textnormal{d}}t}=-\nabla\mathcal{L}(\theta(t)),\\ \text{\ starting with\ }\theta(0)=(\sigma_{\rm init},\cdots,\sigma_{\rm init})^{\top},\end{gathered}$ (12)
@@
According to (9), this lemma implies that, at the end of Training Stage I, the first layer captures the preceding token $x_{s-1}$ for each token $x_{s}$, i.e., $y_{s}=x_{s-1}$. This property is crucial for transformers to implement induction heads and aligns with our approximation result in Theorem 3.1. The proof of Lemma 4.3 is deferred to Appendix D.1.
@@ 4.2 Training Stage II: Transition from $4$-gram to Induction Head
In Training Stage II, it holds that $|{w_{Q}^{(2,2)}}(t)|^{2}\equiv|{w_{K}^{(2,2)}}(t)|^{2}$.
Lemma 4.4 is similar to the balance result for homogeneous networks (Du et al., 2018), and its proof can be found at the start of Appendix D.2. By this lemma, we can define $w_{KQ}^{(2,2)}:=w_{Q}\equiv w_{K}$. Additionally, Lemma 4.3 ensures that $p^{(1,1)}=+\infty$ holds during Stage II. For simplicity, we denote $w_{V_{1}}:=w_{V}^{(2,1)},w_{V_{2}}:=w_{V}^{(2,2)},p:=p^{(2,1)},w_{KQ}:=w_{KQ}^{(2,2)}$. Consequently, the training dynamics are reduced to four parameters
@@
where we still denote the set of parameters as $\theta$ without introducing ambiguity. It is important to note that the problem remains highly non-convex due to the joint optimization of both inner parameters ($p,w_{KQ}$) and outer parameters ($w_{V_{1}},w_{V_{2}}$) in the two heads. At this training stage, GF has a unique fixed point:
$w_{V_{1}}=\frac{\alpha^{\star}}{1+\alpha^{\star}},\,w_{V_{2}}=\frac{1}{1+\alpha^{\star}},\,{p}=+\infty,\,w_{KQ}=w^{\star},$
which corresponds to a global minimizer of the objective (11).
Figure 2: Visualization of the dynamical behavior of Training Stage II with total loss, partial loss, and the parameter evolution. Here, $\alpha^{\star}=1,w^{\star}=0.49,\sigma_{\rm init}=0.01,L=40$. The is clearly shown that transformer learns the $4$-gram component first and then, starts to learn the induction head mechanism. Notably, the entire dynamics unfold in four distinct phases, consistent with our theoretical results (Theorem 4.5). For more experimental details, we refer to Appendix B.3.
As shown in Figure 2, a learning transition from the 4-gram mechanism to the induction head mechanism does occur in our setting. Moreover, the learning process exhibits a four-phase dynamics. The next theorem provides a precise characterization of the four phases, whose proof can be found in Appendix D.2.
@@
    $\mathcal{L}_{\textnormal{{G}}_{4}}(\theta(T_{\rm I}))\leq$ $0.01\cdot\mathcal{L}_{\textnormal{{G}}_{4}}(\theta(0)),$
    $\mathcal{L}_{\textnormal{{IH}}_{2}}(\theta(T_{\rm I}))\geq$ $0.99\cdot\mathcal{L}_{\textnormal{{IH}}_{2}}(\theta(0)).$
@@
    $=$ $\Theta\left((\alpha^{\star}+1)^{2}L\log(1/\sigma_{\rm init})/{w^{\star}}^{2}\right);$
    $T_{\rm III}:=$ $\inf\Big{\{}t>T_{o}:\mathcal{L}_{\textnormal{{IH}}_{2}}(\theta(t))\leq 0.01\cdot\mathcal{L}_{\textnormal{{IH}}_{2}}(\theta(T_{o}))\Big{\}}$
    $=$ $\Theta\left((\alpha^{\star}+1)^{2}L\log(1/\sigma_{\rm init})/{w^{\star}}^{2}\right).$
@@
    $w_{KQ}(t)=\sigma_{{\rm init}}\exp\left(\Theta\left(\frac{{w^{\star}}^{2}t}{(1+\alpha^{\star})^{2}L}\right)\right),\ t<T_{\rm III}.$
  • Phase IV (convergence). In this phase, the loss converges toward zero. Specifically, the following convergence rates hold for all $t>T_{\rm III}:$
    ${\mathcal{L}_{\textnormal{{G}}_{4}}(\theta(t))}$ $=\mathcal{O}\left(\frac{1}{t}\right),$
    $\mathcal{L}_{\textnormal{{IH}}_{2}}(\theta(t))$ $=\mathcal{O}\left(\exp\left(-\Omega\left(\frac{{w^{\star}}^{2}t}{(1+\alpha^{\star})^{2}L}\right)\right)\right),$
@@
Proof idea. We highlight that our fine-grained analysis of entire learning process is guided by two key observations: 1) the dynamics of the two heads can be decoupled; 2) there exist a distinct transition point in the dynamics of each head, as shown in Figure 2 (right). These insights lead us to divide the analysis of each head into two phases: a monotonic phase and a convergence phase. Particularly, for the convergence phase, we introduce a novel Lyapunov function that leverages the unique dynamical structure of self-attention. This Lyapunov function may be of independent interest and offers potential for studying broader issues in self-attention dynamics.
Remark 4.6.
We conduct additional experiments to validate our theoretical insights into the training dynamics and learning transition across a wider range of scenarios. These includes using data distribution (Figure 6) and optimization algorithms (Figure 7) in high-dimensional settings, as well as training real-world transformers on natural language datasets (Figure 5).
5 Experimental Validation
Our theoretical analysis provide insights into the approximation power and optimization dynamics of transformers. To further validate these insights, we conduct a series of experiments, ranging from simple toy models to real-world natural language training tasks. Due to space limitations, the detailed experimental results and setups are provided in Appendix B.
@@ 6 Conclusion
Acknowledgments
Mingze Wang is supported by Young Scientists (PhD) Fund of the National Natural Science Foundation of China (No. 124B2028) and the National Key Basic Research Program of China (No. 2015CB8560). Lei Wu is supported by the National Key R&D Program of China (No. 2022YFA1008200) and National Natural Science Foundation of China (No. 12288101). We thank Dr. Hongkang Yang for insightful discussions and Mingyu Xu, as well as the anonymous reviewers, for their valuable suggestions.
@@ Acknowledgments
Appendix A Related Works
Expressiveness of transformers. Theoretically, Dehghani et al. (2019); Pérez et al. (2021); Wei et al. (2022) explored the Turing-completeness of transformers; Yun et al. (2019) established the universal approximation property of transformers. Subsequent studies examined the efficiency of transformers in representing specific functions or tasks, such as sparse functions (Edelman et al., 2022), targets with nonlinear temporal kernels (Jiang and Li, 2023), practical computer programs (Giannou et al., 2023), long but sparse memories (Wang et al., 2024), induction head (Sanford et al., 2024a; b; Rajaraman et al., 2024), and memorization and reasoning (Chen and Zou, 2024). Besides, many studies suggest that transformers achieve in-context learning by approximating gradient-based iterations across various layers (Garg et al., 2022; Akyürek et al., 2022; Von Oswald et al., 2023; Mahankali et al., 2023; Bai et al., 2023; Shen et al., 2023). Besides, several studies explored the limitation of transformer’s expressivity, particularly in modeling formal languages or simulating circuits (Hahn, 2020; Weiss et al., 2021; Bhattamishra et al., 2020; Merrill et al., 2022; Merrill and Sabharwal, 2023). Among all these works, the most closely related to ours are Rajaraman et al. (2024), which examined a generalized induction head similar to our Eq. (6). Specifically, they showed that multi-layer transformers with single-head attention can implement this mechanism. In contrast, we prove that two-layer transformers are sufficient if multihead attention is used.
Training dynamics of transformers. To gain insights into the dynamics of training transformers, several studies have analyzed simplified transformers on toy tasks. These tasks include learning distinct/common tokens (Tian et al., 2023), leaning balance/inblanced features (Huang et al., 2023), linear regression task (Zhang et al., 2023; Ahn et al., 2024), multi-task linear regression (Chen et al., 2024a), binary classification (Li et al., 2024), transformer with diagonal weights (Abbe et al., 2024), learning causal structure (Nichani et al., 2024), sparse token selection task (Wang et al., 2024), and learning $n$-gram Markov chain (Chen et al., 2024b). Additionally, studies such as those by Ataee Tarzanagh et al. (2023), Tarzanagh et al. (2023) and Vasudeva et al. (2024) have analyzed scenarios where transformers converge to max-margin solutions. Furthermore, Thrampoulidis (2024) has examined the implicit bias of next-token prediction. Among these works, the most closely related to ours are Nichani et al. (2024) and Chen et al. (2024b), which proved that two-layer transformers can converge to induction head solutions. In this work, we explore a setting where the target is a mixture of 4-gram and induction head. We show that two-layer transformers can effectively converge to this mixed target and provide a precise description of the learning process associated with each component. Importantly, we are able to capture the abrupt transition from learning 4-gram patterns to mastering the induction head mechanism—a critical phase in the learning of induction heads, as highlighted in the seminal works (Elhage et al., 2021; Olsson et al., 2022).
Now we discuss the relationship between our work and two closely related studies  (Bietti et al., 2024; Edelman et al., 2024).
Comparison with Bietti et al. (2024).
  • Approximation analysis:
    Bietti et al. (2024) focus primarily on the implementation of the vanilla induction head. In contrast, our study extends this analysis by investigating not only how two-layer transformers achieve vanilla induction heads (Eq. (4)) but also how they implement generalized induction heads, i.e., in-context n-grams (Eqs. (6) and (7)).
    Furthermore, our work provides explicit approximation rate results, offering insights into the distinct roles of multiple heads, positional encoding, dot-product structure, and FFNs in implementing these induction heads.
  • Optimization analysis:
    Study objective: While Bietti et al. (2024) examines the transition from 2-gram to induction head, our work focuses on the transition from 4-gram to induction head.
    study methods: Bietti et al. (2024) conducts extensive experiments supported by partial theoretical properties but does not fully characterize the training dynamics theoretically. In contrast, our study provides a precise theoretical analysis of the entire training process in a toy model, uncovering the sharp transition from 4-gram to induction head.
    Main insights: Bietti et al. (2024) emphasizes the the role of weight matrices as associative memories and the impact of data distributional properties. Our analysis, on the other hand, identifies two primary drivers of the transition: (1) the time-scale separation due to low- and high-order parameter dependencies in self-attention; (2) the speed differences caused by the relative proportions of the two components in the mixed target.
Comparison with Edelman et al. (2024). The primary connection between Edelman et al. (2024) and our work lies in the optimization analysis. Specifically, Edelman et al. (2024) focuses on the transition from uni-gram to bi-gram mechanisms in Markov Chain data. In contrast, our study investigates the transition from 4-gram to in-context 2-gram mechanisms (induction head). Additionally, we theoretically identify two primary drivers of the transition: (1) the time-scale separation due to low- and high-order parameter dependencies in self-attention; (2) the speed differences caused by the relative proportions of the two components in the mixed target.
Appendix B Experiments
B.1 Experiments supporting approximation results
1. Supporting our construction in Theorem 3.3.
Setup. We linear probing experiments (Alain and Bengio, 2016) on the transformers with $H=D=8$ trained in the above experiment (Figure 4). For each checkpoint model ${\rm TF}$, we denote its output in the first layer on the input sequence $X$ as ${\rm TF}^{(1)}(X)$. The probing loss is measured by $\text{dist}\left(X_{\cdot-n+1:\cdot};{\rm TF}^{(1)}(X)\right)=\min\limits_{P\in\mathbb{R}^{D\times n}}:\sum\limits_{s=n}^{L}\left\|X_{s-n+1:s}-{\rm TF}_{s}^{(1)}(X)P\right\|$, where $n=4$, $L=10$, and $X=(x_{1},\cdots,x_{L})$ is generated by $x_{i}\overset{i.i.d.}{\sim}{\rm Unif}(\{\pm 1\})$ with testing batch $1000$. The results are shown in Figure 3.
Figure 3: Probing results supporting our construction in Theorem 3.3. First, we train a two-layer two-layer transformer with head $H=8$ and embedding dimension $D=8$ to learn Eq. (6) with $n=4$, and the checkpoints are stored during training. For each checkpoint model ${\rm TF}$, we denote its output in the first layer on the input sequence $X$ as ${\rm TF}^{(1)}(X)$. To validate whether it encodes the semantic information $X_{s-n+2:s}$ near each $x_{s}$, as predicted by our construction, we conduct a standard linear probing experiment (Alain and Bengio, 2016). Specifically, we measured $\text{dist}\left(X_{\cdot-n+1:\cdot};{\rm TF}^{(1)}(X)\right)=\min\limits_{P\in\mathbb{R}^{D\times n}}:\sum\limits_{s=n}^{L}\left\|X_{s-n+1:s}-{\rm TF}_{s}^{(1)}(X)P\right\|$. As the results shown, the probing loss decreases significantly during training, confirming our key construction in Theorem 4.3: the first layer is responsible for extracting local semantic information $X_{s-n+2:s}$ near each $x_{s}$, enabling the second layer to generate the final output.
2. Supporting the necessity of the required $H$ and $D$ in Theorem 3.3.
Setup. We train two-layer transformers (without FFN layers) with varying $H$ and $D$ to learn the generalized induction head (6) with $n=4$. The input sequence $X=(x_{1},\cdots,x_{L})$ is boolean, with $x_{i}\overset{i.i.d.}{\sim}{\rm Unif}(\{\pm 1\})$ and $L=10$. Each model is trained for 200,000 iterations using squared loss and (online) Adam optimizer with learning rate 5e-4 and batch size $B=100$. Both layers are trained simultaneously. The results for the models with $D=H=8$ and $D=H=2$ are presented in Figure 5.
Figure 4: Results supporting the necessity of the required number of heads $H$ and embedding dimension $D$ in Theorem 3.3. We train two-layer transformers with varying $H$ and $D$ to learn the target in Eq. (6) with $n=4$. The results indicate that the transformer with $H=D=8$ ($>n$) successfully expresses this task, while the transformer with $H=D=2$ ($<n$) fails. These results confirm that the sufficient conditions provided in Theorem 3.3 ($H\gtrsim n$ and $D\geq nd$, where $d=1$ in our setting) are also nearly necessary.
B.2 Additional experiments supporting optimization dynamics
1. Standard transformers on real-world natural language dataset.
Setup. We train a two-layer two-head standard transformer with RPE (3) (without any simplification) on the wikitext-2 dataset, a natural language dataset (Merity et al., 2016). The transformer has an embedding dimension $D=128$ and FFN width $W=512$. For this dataset, the input dimension is $d=33278$. We use a context length $L=200$ and batch size $B=32$. The parameters are initialized with the scale $0.01$. The model is trained for 1,500 epochs on 1 H100, using cross-entropy loss and SGD with learning rate $0.1$, and the initialization scale is $0.01$. It is important to note that both layers are trained simultaneously. The results are presented in Figure 5.
Figure 5: The loss and parameters for the experiment training a two-layer two-head standard transformer (without any simplification) on the wikitext-2 dataset (Merity et al., 2016). Here, $\|p\|$ and $\|(W_{K},W_{Q})\|$ denote the Frobenius norms of all positional encoding parameters and all $W_{K},W_{Q}$ parameters across layers and heads, respectively, The results show that: the loss exhibits a clear plateau; position encoding $p$’s are learned first; and the dot-product structure $W_{K},W_{Q}$ are learned slowly at the beginning, resembling an exponential increase; additionally, as $W_{K},W_{Q}$ are learned, the loss escapes that plateau. These findings closely resemble the behavior observed in our toy model (Figure 2). This experiment provides further support for our theoretical insights regarding the time-scale separation between the learning of positional encoding and the dot-product structure.
2. Discrete token distribution in toy setting.
Setup. We modified the Gaussian input distribution used in the setup for Figure 2 to a boolean input distribution, where each input token, where each input token $x_{i}\stackrel{{\scriptstyle iid}}{{\sim}}{\rm Unif}(\{\pm 1\})$ for $i\in[L]$, All other experimental setups remain the same as in the setup for Figure 2. The training dynamics of Stage (ii) are presented in Figure 6. We can see clearly that the dynamical behavior of the learning process is nearly the same as the one observed for Gaussian inputs in Figure 2.
Figure 6: Visualization of the total loss, partial loss, and the parameter dynamics, for the experiment on discrete token distribution (Boolean, $X\sim{\rm Unif}(\{\pm 1\}^{L})$) in our toy setting with $\alpha^{\star}=1,w^{\star}=0.49,\sigma_{\rm init}=0.01,L=40$. The figure clearly shows that transformer learns the $4$-gram component first and then, starts to learn the induction head mechanism. Notably, the entire dynamics exhibit four phases. These results are extremely similar to that observed with Gaussian inputs, as shown in Figure 2.
3. Adam in high-dimensional toy setting.
Setup. We modified the setup for Figure 2 to employ a high-dimensional model $(D=100)$. Specifically, the target is $w^{\star}=0.49I_{D}/D$, the dot-produce parameters are $W_{K},W_{Q}\in\mathbb{R}^{D}$, initialized such that $\|W_{K}\|_{F},\|W_{Q}\|_{F}=\sigma_{\rm init}$. Additionally, for the Adam optimizer, we use learning rate 5e-4. All other experimental setups remain the same as in the setup for Figure 2.
The training dynamics are depicted in Figure 7, where, for comparison, results using GD are also presented. In both scenarios, the learning process begins with the 4-gram pattern, followed by a gradual learning phase of the induction head mechanism. Notably, within the given number of iterations, GD remains stuck in the plateau, whereas Adam successfully escapes that plateau.
Figure 7: Partial loss for the experiment comparing GD v.s. Adam optimizer in high-dimensional settings ($D=100$). In this setting, a larger $D$ increases the difficulty of the transition from the lazy regime (learning $4$-gram) to the rich regime (learning induction head). The results indicate that: (1) GD learns the 4-gram component first but becomes stuck in a plateau when learning induction head; (2) Adam, while eventually transitioning from the lazy regime (learning 4-gram) to the rich regime (learning induction head), experiences a challenging transition characterized by multiple plateaus during learning induction heads. This finding closely resembles the dynamics for GD.
B.3 Experimental details for Figure 2
In line with our theoretical setting, we examine a simplified two-layer transformer, as described in (10). Specifically, the first layer only contains RPE (3) and the second layer consists of two heads: one uses only RPE and the other employs only dot-product structure. The target function is specified by (8) with $\alpha^{\star}=1,w^{\star}=0.49,\sigma_{\rm init}=0.01,L=40$, and the distribution of each token is Gaussian, i.e., $x_{i}\stackrel{{\scriptstyle iid}}{{\sim}}\mathcal{N}(0,1)$ for $i\in[L]$. Training is conducted by minimizing the squared loss (11) using online SGD with learning rate $0.1$ and batch size $B=1,000$. Following our theoretical analysis, the two layers are trained sequentially:
  • Training Stage I: only the first layer is trained for 100,000 iterations;
  • Training Stage II: Subsequently, only the second layer undergoes training for another 100,000 iterations.
The dynamical behavior of the Training Stage II is visualized in Figure 2.
Appendix C Proofs in Section 3
C.1 Proof of Theorem 3.1
$\textnormal{{IH}}_{2}(X_{L})=(x_{s})_{s=2}^{L-1}\ {{\rm softmax}}\Big{(}\big{(}x_{L}^{\top}W^{\star}x_{s-1}\big{)}_{s=2}^{L-1}\Big{)}^{\top},$ (13)
Theorem C.1 (Restatement of Theorem 3.1).